The authorized share capital of the Alfred Cake Company is 120,000 shares. The equity is currently shown in the company’s books as follows:

How many shares a. issued, b. outstanding and c. how many more shares can be issued without the approval of shareholders?

Common stock ($2 par value) $ 80,000
Additional paid-in capital 30,000
Retained earnings 50,000
Common equity $ 160,000
Treasury stock (2,000 shares) 24,000
Net common equity $ 136,000
Answer #1

a) Issued shares = Common stock / Par value = 80,000 / 2 = 40,000

b) Treasury shares = 24,000 / 2 = 12,000

Outstanding shares = 40,000 - 12,000 = 28,000

c) More shares that can be issued = 120,000 - 28,000 = 92,000